OpenStudy (anonymous):

A ladder is propped up 3 feet away from the side of a building. The length of the ladder is 8 feet. What is the height of the building? Round to the nearest tenth. A. 8.5 ft B. 7.4 ft C. 7.6 ft D. 9 ft

OpenStudy (across):

Wow... those are very short buildings...

OpenStudy (anonymous):

lol, right

OpenStudy (anonymous):

i thought it would be C but i need a 2nd opinion

OpenStudy (across):

Use the Pythagorean theorem\[a^2+b^2=c^2,\]with \(a=3\) and \(c=8\).

OpenStudy (anonymous):